OpenStudy (anonymous):

I will give a MEDAL if you get this right! Find an equation of the circle that has center C(8, −5) and is tangent to the line y = 5. (Give your answer in standard notation.)

OpenStudy (anonymous):

formula for the eq of circle at C(8,-5)=(h,k) and tangent yo the line y=5=radius r just plug into the formula (x-h)^2+(y-k)^2=r^2

OpenStudy (anonymous):

yeah i got (x-8)^2+(y+5)^2=5^2 and its wrong

OpenStudy (anonymous):

that was correct my be you just need to simplified more like (x-8)^2+(y+5)^2=5^2 (x-8)^2+(y+5)^2=25 or simplify more